93 passengers rode in a train. Tickets for regular coach seats cost $124. Tickets for sleeper cars seats cost $292. The receipts for the trip totaled $20604. How many passengers purchased each type of ticket? Coach? Sleeper car?
Answer by addingup(3677)

coach : c
sleeper: s
~~~~~~~~~~~~~~~~~~~
c+s = 93 subtract c from both sides:
s = 93 - c
~~~~~~~~~~~~~~~~~~~~~~~~~~~~
124c + 292s = 20604
substitute for s:
124c + 292(93 - c) = 20604
124c + 27156 - 292c = 20604
-168c = -6552 divide, and remember that -/- = +
c = 39
and s = 93 - c = 93 - 39 = 54